The width of a rectangle is one-half of the length. The perimeter of the rectangle is 90. What are the dimensions of the rectangle?
Answer by checkley77(12569)

W=L/2
2(W+L)=90
W+L=90/2
W+L=45
L/2+L=45
(L+2L)/2=45
3L/2=45
3L=90
L=90/3
L=30 ANS. FOR THE LENGTH
W=30/2
W=15 ANS. FOR THE WIDTH.
pROOF:
2(30+15)=90
2*45=90
90=90